OpenStudy (anonymous):

Solve the equation. (x^2 - 3)(x^2 + 3x - 4) = 0

OpenStudy (cwrw238):

either x^2 - 3 = 0 or x^2 = 3x - 4 = 0 solve these 2 equations you should find 4 roots